OpenStudy (anonymous):

Sketch the region enclosed by y=e^{2 x}, y=e^{7 x}, and x=1. Decide whether to integrate with respect to x or y. Then find the area of the region.

OpenStudy (anonymous):

\[y=e ^{2x} y=e ^{7x} x=1\]

OpenStudy (anonymous):

hmm did you sketch it already?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Where in the heck have you been?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

My pup wanted out... wanted to eat a squirrel.

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

cute

OpenStudy (anonymous):

Testing 123

OpenStudy (dumbcow):

for a bounded region \[Area = \int\limits_{a}^{b}f(x) - g(x) dx\] where f(x) > g(x) in interval [a,b]

OpenStudy (anonymous):

Yea you lost me

OpenStudy (anonymous):

Teach it to me like Im a 2 year old

OpenStudy (anonymous):

haha

OpenStudy (dumbcow):

haha i know its like i just gave you a theorem out of a textbook ... anyway, you have 2 functions right...thats your f(x) and g(x) graph them and see which one is above the other, also find where they intersect e^2x = e^7x

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\[\int\limits_{0}^{1} (e ^{7x} -e ^{2x}) \]

OpenStudy (anonymous):

ANd we got this from

OpenStudy (anonymous):

I mean ho

OpenStudy (anonymous):

how

OpenStudy (anonymous):

OH wait is the area when I plug it in?

OpenStudy (anonymous):

top-bottom from 0 to 1...if you integrate with respect to x

OpenStudy (dumbcow):

he got it from graphing and solving for where they intersect :)

OpenStudy (anonymous):

graph.... do you have a graphing calculator?

OpenStudy (anonymous):

Yea but I never used it haha

OpenStudy (anonymous):

lol... you have to use it... or algebra knowledge

OpenStudy (anonymous):